Calculate the kinetic energy (inMeV) of an alpha particle whose DeBrogile wavelength is equal to 1 femtometer (1E-15m), which is about the size of a nucleus. The mass of the alpha particle is 6.65E-27kg.

answer choices:

3.3E-11 MeV

5.3E-30MeV

49.8MeV

103MeV

206MeV
Answer #1

(6.626 X16 -०५ Js) >15 6.626 X10 x 10-19 Heg-mols IX1015 þ = me 2 6626x10 6065x1627 /s 2. = 9.964x107 mie KE = olue -1) 3030